# Interactive Sketchpad: An Interactive Multimodal System for Collaborative, Visual Problem-Solving

### v1

@@ Abstract.
Humans have long relied on visual aids like sketches and diagrams to support reasoning and problem-solving. Visual tools, like auxiliary lines in geometry or graphs in calculus, are essential for understanding complex ideas. However, many tutoring systems remain text-based, providing feedback only through natural language. Leveraging recent advances in Large Multimodal Models (LMMs), this paper introduces Interactive Sketchpad, a tutoring system that combines language-based explanations with interactive visualizations to enhance learning. Built on a pre-trained LMM, Interactive Sketchpad is fine-tuned to provide step-by-step guidance in both text and visuals, enabling natural multimodal interaction with the student. Accurate and robust diagrams are generated by incorporating code execution into the reasoning process. User studies conducted on math problems such as geometry, calculus, and trigonometry demonstrate that Interactive Sketchpad leads to improved task comprehension, problem-solving accuracy, and engagement levels, highlighting its potential for transforming educational technologies.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXccs: Human-centered computing Human computer interaction (HCI)$c$s: Human-centered computing Interaction techniquesccs: Computing methodologies Machine learningccs: Computing methodologies Artificial intelligence
@@ 3. Interactive Sketchpad
1) Problem Analysis. We prompt a pre-trained GPT-4o model (OpenAI, 2024) regarding when visualizations may be helpful based on the problem and user query (see examples in Appendix D.1).
@@
3) Hint Generation. We prompt a GPT-4o model to give subtle hints rather than revealing the answer and to generate visualizations for the hint using code. To enable good hints and appropriate visual hints, we collected human demonstrations of good hints interleaved with images in an example conversation between a student and a tutor, using several samples of math and coding problems. We then fine-tuned the model with this data (prompt found in Appendix D.1). In subsequent iterations, the model uses the conversation history, a screenshot of the user’s interactive whiteboard, and the user’s query to generate hints. Finally, we also provide the model access to the OpenAI Code Interpreter to perform mathematical calculations by executing code to reduce calculation errors, similar to Chen et al. (2023). For example, the model will perform calculations such as numerical integration for calculus questions, or arithmetic for systems of equations to provide more accurate hints and check student answers (see examples in Appendix C).
@@ 4.2. Performance Comparisons
We first run experiments evaluating the performance of Interactive Sketchpad on directly solving the problem without giving hints to evaluate whether the system can solve the problem itself. We observe in Table 1 that Interactive Sketchpad consistently outperforms Visual Sketchpad on IsoBench. Visual Sketchpad’s prompts encourage the LMM to use its own visual reasoning to solve problems rather than use code execution to help solve the problems. Interactive Sketchpad on the other hand does not specify the same constraint to not use code execution, instead allowing its use to improve problem-solving accuracy. This highlights the effectiveness of combining both code execution for calculations and visual reasoning in solving complex problems. We remove the constraint of using visual reasoning rather than code execution as we want to make the system achieve as high accuracy as possible to reduce errors that may confuse the student. We use the original IsoBench prompts (Fu et al., 2024), appending a prompt footer (see D.4) that instructs the model to explain its answer and format the output accordingly. Additionally, we employ a system prompt (see D.3) directing Interactive Sketchpad to solve the question directly.
@@ 4.3. User Studies
The primary goal of the user study was to assess the system’s ability to enhance the learning experience by fostering active engagement and improving conceptual understanding with visual artifacts. (Shatri and Buza, 2017) The selected problems mirror the type of challenges that benefit from visual support in mathematic context, such as graphs and functions. (Refer to the Appendix A.1). Since one of the core features of Interactive Sketchpad is its integration of visual and textual feedback, a key consideration for our user study was to analyze the quality of the system’s multimodal interaction. To do so, we measured how students engaged with these modalities, focusing on the effectiveness of visual explanations in complementing or enhancing textual guidance. Another key consideration was the evaluation of our system’s iterative learning process. To evaluate this, students were encouraged to iterate through problems to explore how the feedback influenced their learning experience.
@@ Appendix A Appendix
A.1. User Study Details
To ensure that our findings could translate to practical use cases, we tested with mathematical problems that could efficiently benefit from visual artifacts. Questions added to the users’ question bank were sourced from previous exam questions of the Scholastic Aptitude Test (SAT), the IsoBench dataset (Fu et al., 2024), a previous study on visualization for math (Rösken and Rolka, 2006) and Geometry3k (Lu et al., 2021). Visualization plays an important role when learning mathematics, and a good example is solving an integral problem. Visual interpretations of the function and shading parts of the graph can significantly benefit the students’ learning experience, which has been demonstrated in previous studies (Rösken and Rolka, 2006). By selecting both practical mathematical topics that range across various difficult levels, we ensured the system catered to a broad range of learners. The system’s potential for versatile use across subjects and learning environments allows our results to demonstrate foundational interaction between AI and humans for education.
Appendix B User Interface
Interactive Whiteboard Interface: When the user wants to draw, write math or annotate on the diagram given by Interactive Sketchpad , they can do so on the interactive whiteboard a natural way to do this is by using an iPad and Apple Pencil. They can then share the working on the interactive whiteboard with Interactive Sketchpad using the ‘Send Screenshot‘ button. Visualizations generated by Interactive Sketchpad will automatically appear as a new page on the interactive whiteboard and users can toggle between pages to access their working at various stages of solving different problems.
Refer to caption Refer to caption Refer to caption
Appendix C Mathematical Calculations via Code Execution
@@
Appendix D Prompts
D.1. System Prompt for Hint Generation & Visualization Generation using Code
@@ 2. You should be concise with your response
D.3. System Prompt for Direct Problem Solving
@@
D.4. Prompt Footer for IsoBench

### v2

# Interactive Sketchpad: A Multimodal Tutoring System for Collaborative, Visual Problem-Solving

@@ Abstract.
Humans have long relied on visual aids like sketches and diagrams to support reasoning and problem-solving. Visual tools, like auxiliary lines in geometry or graphs in calculus, are essential for understanding complex ideas. However, many tutoring systems remain text-based, providing feedback only through natural language. Leveraging recent advances in Large Multimodal Models (LMMs), this paper introduces Interactive Sketchpad, a tutoring system that combines language-based explanations with interactive visualizations to enhance learning. Built on a pre-trained LMM, Interactive Sketchpad is fine-tuned to provide step-by-step guidance in both text and visuals, enabling natural multimodal interaction with the student. Accurate and robust diagrams are generated by incorporating code execution into the reasoning process. User studies conducted on math problems such as geometry, calculus, and trigonometry demonstrate that Interactive Sketchpad leads to improved task comprehension, problem-solving accuracy, and engagement levels, highlighting its potential for transforming educational technologies. All code is available at: https://stevenshinechen.github.io/interactivesketchpad/.
@@
copyright: acmlicensedjournalyear: 2025doi: 3706599.3719790journalyear: 2025copyright: rightsretainedconference: Extended Abstracts of the CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: Extended Abstracts of the CHI Conference on Human Factors in Computing Systems (CHI EA ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706599.3719790isbn: 979-8-4007-1395-8/2025/04ccs: Human-centered computing Human computer interaction (HCI)$c$s: Human-centered computing Interaction techniquesccs: Computing methodologies Machine learningccs: Computing methodologies Artificial intelligence
@@
[…]
This image shows an example overview of how a user can interact with Interactive Sketchpad
@@
[…]
This image shows a detailed example overview of how a user can interact with Interactive Sketchpad, including how code generation is utilized to create textual and visual hints.
@@ 2. Related Work
[…]
This image shows a flowchart of the system flow, along with a screenshot example of a whiteboard interaction.
@@ 3. Interactive Sketchpad
1) Problem Analysis. We prompt a pre-trained GPT-4o model (OpenAI, 2024) regarding when visualizations may be helpful based on the problem and user query (see examples in Appendix B.1).
@@
3) Hint Generation. We prompt a GPT-4o model to give subtle hints rather than revealing the answer and to generate visualizations for the hint using code. To enable good hints and appropriate visual hints, we collected human demonstrations of good hints interleaved with images in an example conversation between a student and a tutor, using several samples of math and coding problems. We then fine-tuned the model with this data (prompt found in Appendix B.1). In subsequent iterations, the model uses the conversation history, a screenshot of the user’s interactive whiteboard, and the user’s query to generate hints. Finally, we also provide the model access to the OpenAI Code Interpreter to perform mathematical calculations by executing code to reduce calculation errors, similar to Chen et al. (2023). For example, the model will perform calculations such as numerical integration for calculus questions, or arithmetic for systems of equations to provide more accurate hints and check student answers (see examples in Appendix A.1).
@@
3.1. User Interface
[…]
Interactive Whiteboard Interface: When the user wants to draw, write math or annotate on the diagram given by Interactive Sketchpad , they can do so on the interactive whiteboard. A natural way to do this is by using an iPad and Apple Pencil. They can then share the working on the interactive whiteboard with Interactive Sketchpad using the ‘Send Screenshot‘ button. Visualizations generated by Interactive Sketchpad will automatically appear as a new page on the interactive whiteboard and users can toggle between pages to access their working at various stages of solving different problems.
@@ 4.2. Performance Comparisons
We first run experiments evaluating the performance of Interactive Sketchpad on directly solving the problem without giving hints to evaluate whether the system can solve the problem itself. We observe in Table 1 that Interactive Sketchpad consistently outperforms Visual Sketchpad on IsoBench. Visual Sketchpad’s prompts encourage the LMM to use its own visual reasoning to solve problems rather than use code execution to help solve the problems. Interactive Sketchpad on the other hand does not specify the same constraint to not use code execution, instead allowing its use to improve problem-solving accuracy. This highlights the effectiveness of combining both code execution for calculations and visual reasoning in solving complex problems. We remove the constraint of using visual reasoning rather than code execution as we want to make the system achieve as high accuracy as possible to reduce errors that may confuse the student. We use the original IsoBench prompts (Fu et al., 2024), appending a prompt footer (see B.4) that instructs the model to explain its answer and format the output accordingly. Additionally, we employ a system prompt (see B.3) directing Interactive Sketchpad to solve the question directly.
@@ 4.3. User Studies
The primary goal of the user study was to assess the system’s ability to enhance the learning experience by fostering active engagement and improving conceptual understanding with visual artifacts. (Shatri and Buza, 2017) To ensure that our findings could translate to practical use cases, we tested with mathematical problems that could efficiently benefit from visual artifacts. Questions added to the users’ question bank were sourced from previous exam questions of the Scholastic Aptitude Test (SAT), the IsoBench dataset (Fu et al., 2024), a previous study on visualization for math (Rösken and Rolka, 2006) and Geometry3k (Lu et al., 2021). Visualization plays an important role when learning mathematics, and a good example is solving an integral problem. Visual interpretations of the function and shading parts of the graph can significantly benefit the students’ learning experience, which has been demonstrated in previous studies (Rösken and Rolka, 2006). By selecting both practical mathematical topics that range across various difficult levels, we ensured the system catered to a broad range of learners. The system’s potential for versatile use across subjects and learning environments allows our results to demonstrate foundational interaction between AI and humans for education.
Since one of the core features of Interactive Sketchpad is its integration of visual and textual feedback, a key consideration for our user study was to analyze the quality of the system’s multimodal interaction. To do so, we measured how students engaged with these modalities, focusing on the effectiveness of visual explanations in complementing or enhancing textual guidance. Another key consideration was the evaluation of our system’s iterative learning process. To evaluate this, students were encouraged to iterate through problems to explore how the feedback influenced their learning experience.
4.3.3. Qualitative Results.
@@ Appendix A Appendix
A.1. Mathematical Calculations via Code Execution
@@
A.2. User Study Questionnaire
Refer to caption Refer to caption Refer to caption
Appendix B Prompts
B.1. System Prompt for Hint Generation & Visualization Generation using Code
@@ 2. You should be concise with your response
B.3. System Prompt for Direct Problem Solving
@@
B.4. Prompt Footer for IsoBench